Grade 12Algebra

if (a+1)(b+1)(c+1)=4 abc prove that a+b+c=abc
plz help me fast guys exams are near

Ankit Jaiswal
10 Years agoGrade 12

6 Answers

Vikas TU
10 Years ago
Multiply all the three terms and get,
(a+b+c) + (ab + bc +ca) = abc + 2abc
now from property:
(ab +bc +ca) = 2abc
hence,
 a+b+c=abc
Ajay
10 Years ago
How is this statement true? Neither the question nor answer seem to be correct.
now from property:
(ab +bc +ca) = 2abc
Vikas TU
10 Years ago
the qstn. is somewhat partially incompleted.
he should have given one more relation just like similar to what i already meant used.
by my side.
Ankit Jaiswal
10 Years ago
i am sorry about it but i thought it will not help actually in the question it also given that a,b,c are 3 different +ve integers
 
mycroft holmes
10 Years ago
Write the equation as Suppose And so So we have c =1.The equation becomes (1+a)(1+b) = 2ab or 1+a+b = ab.Since a divides ab and a, we have a divides (b+1) and similarly b divides a+1.This yields Also a divides b+1 means a is not equal to b (as b does not divide b+1). So a = b+1.So the equation (a+1)(b+1) = 2ab now becomes a(a+1) = 2a (a-1) so a=3, and b=2.So (a,b,c) = (3,2,1) is the only solution. This solution satisfies the relation a+b+c=abc
mycroft holmes
10 Years ago
Write the equation as 
 
Suppose 
 
And so
 
 
So we have c =1.
 
The equation becomes (1+a)(1+b) = 2ab or 1+a+b = ab.
 
Since a divides ab and a, we have a divides (b+1) and similarly b divides a+1.
 
This yields 
 
Also a divides b+1 means a is not equal to b (as b does not divide b+1). So a = b+1.
 
So the equation (a+1)(b+1) = 2ab now becomes a(a+1) = 2a (a-1) so a=3, and b=2.
 
So (a,b,c) = (3,2,1) is the only solution. This solution satisfies the relation a+b+c=abc